40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog, Page 248 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  5518   Thu Sep 22 13:56:56 2011 kiwamuUpdateASCC1ASS : status update

The output matrix in the C1ASS servo were coarsely readjusted and the servos seemed working.

However it is difficult to say the servo is very good or so-so,

because the ETMY suspension moves a lot and hence the cavity eigen axis moves a lot too.

 


(to do)

 + optimization of the ETMY oplevs and OSEM damping.
 + evaluation of the performance of the C1ASS with a good damping.

(Background)

 Since we have installed the new mid-HV amplifier for the PZT1 mirror (#5450) it changed the response of the PZT1 (gains from EPICS to the actual angles).
Therefore the C1ASS output matrix needs be adapted to the new PZT1 response.
 
(What I have done)
  What I was measuring was a coupling from each PZT mirror to both beam angle and beam position by looking at the output from the LOCKINs.
So this measurement eventually gives us a nicely diagonalized output matrix by inverting the coupling.
However the measurement turned out to be difficult because the ETMY moved too much.
In fact the cavity eigen axis also moves and the fluctuation was larger than the intentionally introduced beam angle/translation offsets, which are for the coupling measurement.
 
 Instead of measuring the couplings, I put some numbers into the matrix based on a guess.
Since the PZT1 HV amp became weaker than that of PZT2, the elements in the output matrix should be amplified by some number.
Right now the PZT1 amp can drive the mirror in a range of -5 -30 V with EPICS range of +/-10 counts, and for PZT2 it is about 0 -150V with EPICS range of +/-5 counts.
So the difference of the responses in unit of V/counts is about 8.5.
The PZT1 elements in the matrix were multiplied by this number and I became able to close the servos.

Quote from #5455

  + Modification of C1ASS (Kiwamu)

  5517   Thu Sep 22 13:45:17 2011 PaulUpdateSUSETMX actuator response fits

Fitting results: 

 Pitch

-- Fit completed after 305 iterations--
 Started with: Gain = 3e-05,
 Q factor = 5,
 Pole frequency = 0.6776,
 Fit results:  Gain = 1.85497e-06,
 Q factor = 23.7233,
 Pole frequency = 0.956686
 Residual (normalised against the sum of input datapoints) = 0.0202483
 
Yaw
-- Fit completed after 334 iterations--
 Started with: Gain = 3e-05,
 Q factor = 5,
 Pole frequency = 0.6776,
 Fit results:  Gain = 2.518e-06,
 Q factor = 7.21618,
 Pole frequency = 0.853559
 Residual (normalised against the sum of input datapoints) = 0.0570132
Attachment 1: ETMX_PITCH_actuator_response_complex_fit.png
ETMX_PITCH_actuator_response_complex_fit.png
Attachment 2: ETMX_YAW_actuator_response_complex_fit.png
ETMX_YAW_actuator_response_complex_fit.png
  5516   Thu Sep 22 11:50:37 2011 KojiUpdateLSCLocking status update

Both loops basically have no phase margins. i.e. unstable. How can you lock PRMI with these servos?

Quote:

The following rough swept sine plots are the OLTs for MICH and PRCL. The gain setting was -10 and 0.5 for MICH and PRCL, respectively. Integrators were off. Looking at the measured plots, MICH has about 300 Hz UGF, when the gain is -20, and PRCL has about 300 HZ UGF, too, when the gain is 0.8.

 

  5515   Thu Sep 22 11:49:05 2011 kiwamuUpdateLSCsome LSC scripts don't run on pianosa

Found some LSC scripts didn't run on pianosa. Particularly all the scripts on the C1:IFO_CONFIGURE screen don't run.

They need to be fixed.

  5514   Thu Sep 22 10:43:50 2011 PaulUpdateSUSPower spectrum with different filter gains

 I thought it might be informative before trying to optimise the filter design to see how the current one performs with different gain settings. I've plotted the power spectra for ITMY yaw with filter gains of 0, 1, 2, 3 and 4.

All of the higher gains seem to perform better than the 0 gain, so can I deduce from this that so far the oplev control loop isn't adding excess noise at these frequencies?

Attachment 1: ITMY_YAW_closed_vs_open_noise.pdf
ITMY_YAW_closed_vs_open_noise.pdf
  5513   Thu Sep 22 04:49:14 2011 AnamariaUpdateLSCLocking status update - Some Scripts, No Louck

The scripts I wrote can be found in /users/anamaria/scripts/sensemat/

]There are two of them:

- one that sets all the switches, gains, frequencies, etc, then cycles through the various RFPDs I and Q into the LOCKIN signal, so as to see the sensing matrix.

- the second one is a matlab script that takes the crappy file tdsavg outputs and makes it into a cute mag/phase matrix.

They're quite primitive at this point, I've forgotten a lot of tcsh... may improve later. But could be useful later to someone else at least.

I don't think it's particularly the fault of the script that we can't measure the sensing matrix. We can slam on the excitation by hand, and it holds for a little while. I set a wait time for lock to adjust, and most times it just oscillates a bit for a few seconds. Also, the script turns on the excitation and it's done, the rest is just measurement, then turns it off at the end. So during the script, there's not much to deal with, except keeping the lowpass filters quiet when switching the signal to demod; but that doesn't go anywhere, so it definitely doesn't disturb the ifo. Turns out pressing the RSET clear history button needs a 2 to make it happen.

I think I might prefer to set the excitation to run, and then do the old retrieve-data-later-nds-matlab thing. I do not trust these measurements without coherence and a bit of variance study, given instabilities.

Point is... Even on carrier, the PRC lock is not stable by any means. Can barely turn on low freq boosts, every other lock. Until we fix the lock stability issue, there's not much to measure I guess.

Unfortunately, I don't know how to make that happen. Before we leave on Friday we could do a few sanity checks such as measuring the noise of the RFPDs vs ADC+whitening, which I may have said I would do; and perhaps setting up a couple OSAs, one on REFL, one on AS, to make sure we know what the sidebands are doing. Both of which Rana suggested at some point.

(There used to be a quote here from Keiko here but I got mad when it reformated my entire log to be one cluster- hence the look)

  5512   Thu Sep 22 01:45:41 2011 KeikoUpdateLSCLocking status update

Keiko, Anamaria

Tonight we want to measure the LSC matrix for PRMI and compare the simulation posted last night (#5495).

First. we locked MICH and PRCL, and measured the OLT to see how good the locking is. The following rough swept sine plots are the OLTs for MICH and PRCL. The gain setting was -10 and 0.5 for MICH and PRCL, respectively. Integrators were off. Looking at the measured plots, MICH has about 300 Hz UGF, when the gain is -20, and PRCL has about 300 HZ UGF, too, when the gain is 0.8.

MICH-OLT.pdf

PRCL-OLT.pdf

As these lokings seemed good, so we tried the LSC matrix code written by Anamaria. However it is not working well at this point. When the script add excitations to the exc channels, they kick the optics too much and the lockings are too much disturbed...

Also, we have been trying to lock PRC with the SB resonant, it doesn't work. Looking at the simulated REFL11I (PRCL) signal (you can see it in #5495 too), the CR and SB resonances have the opposite signs... But minus gain never works for PRCL. It only excites the mirror rather than locking.

  5511   Thu Sep 22 01:05:28 2011 KatrinUpdateGreen LockingNew modulation frequency (Y arm)

[Kiwamu / Katrin]

 

On Wednesday, the green light was locked to the Y arm cavity.

Modulation frequency was changed from 279kHz to 178875Hz. The amplitude was changed from 10Vpp to 0.01Vpp to achieve a modulation index of 0.38. The modulation frequency was changed to minimize AM. With the new modulation frequency the laser light could still locked to the cavity.

The signal of the LO and the photodiode are multiplied by a ZAD-8 mini circuit mixer (Level 7). This mixer requires LO input is +7dBm = 1.4Vpp. Thus we put a 36dB attenuator between the LO and the PZT at the laser. PDH error signal shows lots of peaks that are most likely higher order sidebands. Thus, the next step is to work on the low-pass filter. However the SNR of the error signal has improved with the new modulation frequency. With the old mod. frequency the PDH signal was 4mVpp and the noise floor was 2mVpp.

Phase between the photodiode signal and LO is shifted by about 10 degrees. Step two is to work on a phase shifter.

 

 

  5510   Thu Sep 22 00:00:10 2011 PaulUpdateSUSITMY and SRM actuator response functions - complex fitting results

Here are the results of the complex fitting. The residuals are bigger this time, but still probably small enough to be ok(?), with the possible exception of ITMY PITCH (due again I think to the data points straddling the resonance).

ITMY YAW actuator response complex fit

-- Fit completed after 282 iterations--

 Started with: Gain = 3e-05,
 Q factor = 5,
 Pole frequency = 0.6776,
 Fit results:  Gain = 1.14673e-06,
 Q factor = 12.9471,
 Pole frequency = 0.766531
 Residual (normalised against the sum of input datapoints) = 0.0688174
 
ITMY PITCH actuator response complex fit
-- Fit completed after 191 iterations--
 Started with: Gain = 3e-05,
 Q factor = 5,
 Pole frequency = 0.6776,
 Fit results:  Gain = 1.25105e-06,
 Q factor = 3.88981,
 Pole frequency = 0.706744
 Residual (normalised against the sum of input datapoints) = 0.144165
 
SRM YAW actuator response complex fit
-- Fit completed after 246 iterations--
 Started with: Gain = 3e-05,
 Q factor = 5,
 Pole frequency = 0.6776,
 Fit results:  Gain = 3.34137e-06,
 Q factor = 9.6875,
 Pole frequency = 0.854913
 Residual (normalised against the sum of input datapoints) = 0.0153646
 
SRM PITCH actuator response complex fit
-- Fit completed after 266 iterations--
 Started with: Gain = 3e-05,
 Q factor = 5,
 Pole frequency = 0.6776,
 Fit results:  Gain = 7.97529e-06,
 Q factor = 7.63888,
 Pole frequency = 0.568227
 Residual (normalised against the sum of input datapoints) = 0.0319653
Attachment 1: ITMY_PITCH_actuator_response_complex_FIT.png
ITMY_PITCH_actuator_response_complex_FIT.png
Attachment 2: ITMY_YAW_actuator_response_complex_FIT.png
ITMY_YAW_actuator_response_complex_FIT.png
Attachment 3: SRM_PITCH_actuator_response_complex_FIT.png
SRM_PITCH_actuator_response_complex_FIT.png
Attachment 4: SRM_YAW_actuator_response_complex_FIT.png
SRM_YAW_actuator_response_complex_FIT.png
  5509   Wed Sep 21 23:44:45 2011 PaulUpdateSUSRe:ITMY and SRM actuator response functions - fitting results

Quote:

Did you take the 180 deg shift into your account ?

Since your measurement was done when the loop was closed, there must be an additional 180 deg phase shift (in other words, minus sign).

Quote from #5507

In the end I just fitted the response magnitude. I was initially fitting the complex response function, but ran into problems which I think were cased by overall phase offsets between the data and test function. Can I canvass for opinion if fitting the magnitude is OK, or should I try again fitting the phase too?

 I thought I had, but apparently not, and I'd made another error or two in the complex version of my fitting routine. I've fixed them now, thanks! I'll put up the new fitting results tomorrow morning.

  5508   Wed Sep 21 23:25:51 2011 kiwamuUpdateSUSRe:ITMY and SRM actuator response functions - fitting results

Did you take the 180 deg shift into your account ?

Since your measurement was done when the loop was closed, there must be an additional 180 deg phase shift (in other words, minus sign).

Quote from #5507

In the end I just fitted the response magnitude. I was initially fitting the complex response function, but ran into problems which I think were cased by overall phase offsets between the data and test function. Can I canvass for opinion if fitting the magnitude is OK, or should I try again fitting the phase too?

  5507   Wed Sep 21 23:05:16 2011 PaulUpdateSUSITMY and SRM actuator response functions - fitting results

 I used an fminsearch function to fit the SRM and ITMY actuator response magnitudes. The testfunction was just that for a single second order pole, but it gave what I consider to be good fits for the following reasons:

*for 3 of the 4 fits the residuals were less than 0.5% of the summed input data points. The worst one (ITMY pitch) was about 2.7%, which I think is due to the resonance happening to be right in the middle of two data points.

*the tolerance of 1 part in 10^9 was reached quickly from not very finely tuned starting points.

The test function was: G=abs(Gp./(1+1i.*f./fp./Qp-(f./fp).^2)), where G(f) is the actuator response magnitude, Gp is the pole gain, fp is the pole frequency, and Qp is the pole Q factor.

In the end I just fitted the response magnitude. I was initially fitting the complex response function, but ran into problems which I think were cased by overall phase offsets between the data and test function. Can I canvass for opinion if fitting the magnitude is OK, or should I try again fitting the phase too?

Anyway, here are the results of the fits, and I've attached plots of each too (each one in linear and log y axis because each on its own might be misleading for fits):

EDIT - I added more points to the otherwise sparse looking fitted curves

 

ITMY PITCH actuator response fit

-- Fit completed after 190 iterations--

 Started with: Gain = 3e-06,

 Q factor = 5,

 Pole frequency = 1,

 Fit results:  Gain = 1.32047e-06,

 Q factor = 4.34542,

 Pole frequency = 0.676676

 Residual (normalised against the sum of input datapoints) = 0.0268321

 

ITMY YAW actuator response fit

-- Fit completed after 156 iterations--

 Started with: Gain = 3e-06,

 Q factor = 5,

 Pole frequency = 1,

 Fit results:  Gain = 1.14456e-06,

 Q factor = 8.49875,

 Pole frequency = 0.730028

 Residual (normalised against the sum of input datapoints) = 0.00468077

 

SRM PITCH actuator response fit

 -- Fit completed after 192 iterations--

 Started with: Gain = 3e-06,

 Q factor = 5,

 Pole frequency = 1,

 Fit results:  Gain = 7.94675e-06,

 Q factor = 7.16458,

 Pole frequency = 0.57313

 Residual (normalised against the sum of input datapoints) = 0.00301265

 

SRM YAW actuator response fit

 -- Fit completed after 156 iterations--

 Started with: Gain = 3e-06,

 Q factor = 5,

 Pole frequency = 1,

 Fit results:  Gain = 3.34179e-06,

 Q factor = 9.57601,

 Pole frequency = 0.855322

 Residual (normalised against the sum of input datapoints) = 0.000840468

Attachment 1: ITMY_PITCH_actuator_response_FIT.png
ITMY_PITCH_actuator_response_FIT.png
Attachment 2: ITMY_YAW_actuator_response_FIT.png
ITMY_YAW_actuator_response_FIT.png
Attachment 3: SRM_PITCH_actuator_response_FIT.png
SRM_PITCH_actuator_response_FIT.png
Attachment 4: SRM_YAW_actuator_response_FIT.png
SRM_YAW_actuator_response_FIT.png
  5506   Wed Sep 21 21:13:35 2011 ranaUpdateIOOAM modulation misery

How about changing the x-axis of all these plots into meters or picometers and tell us how wide the PRC resonance is? (something similar to the arm cavity linewidth expression)

Also, there's the question of the relative AM/PM phase. I think you have to try out both I & Q in the sim. I think we expect Q to be the most effected by AM.

  5505   Wed Sep 21 19:20:41 2011 SureshUpdateIOOPSL beam into MC was off in Pitch. Readjusted.

I found the PSL beam into the MC off in pitch by large amount.  I readusted the PSL beam for optimal coupling.

The beam had shifted on the WFS as well.  So I recentered the DC signal on the WFS with the MC unlocked.  However both the DC and RF signals on the WFS shift when we lock the MC.  This ought to indicate sub-optimal coupling of PSL into MC.  But instead, if we were to reduce these offsets on the WFS by adjusting the MC axis it leads to higher reflected power from the MC.

The current plan is to retain these RF offsets and lock the WFS with a DC offset in the servo filters.

  5504   Wed Sep 21 18:53:03 2011 KeikoUpdateIOOAM modulation misery

The signal offset due to the AM modulation is estimated by a simulation for PRCL for now. Please see the result below.

Too see how bad or good the AM modulation with 1/50 modulation depths of PM, I ran a simulation. For example I looked at PRCL sweep signal for each channel. I tried the three AM modulation depths, (1) m_AM=0 & m_PM = 0.17 (2) m_AM = 0.003 & m_PM = 0.17 which is the current modulation situation (3) m_AM = 0.17 & m_PM = 0.17 in which AM is the same modulation depth as PM.  For the current status of (2), there are offsets on signals up to 0.002 while the maximum signal amplitude is 0.15. I can't tell how bad it is.... Any suggestions?

 

(1) m_AM=0 & m_PM = 0.17. There is no offset in the signals.

AM0.png

(2) m_AM = 0.003 & m_PM = 0.17. There are offsets on signals up to 0.002 while the maximum signal amplitude is 0.15.

AMratio50.png

(3) m_AM = 0.17 & m_PM = 0.17. There are offsets on signals up to 0.1 while the maximum signal amplitude is 0.2.

AMratio1.png

I will look at MICH and SRCL in the same way. 

Quote:

I'd like to see some details about how to determine that the ratio of 1:50 is small enough for AM:PM.

* What have people achieved in past according to the elogs©  of the measurements?

* What do we expect the effect of 1:50 to be? How much offset does this make in the MICH/PRC/SRC loops? How much offset is too much?

Recall that we are using frontal modulation with a rather small Schnupp Asymmetry...

 

  5503   Wed Sep 21 17:42:35 2011 ranaUpdateIOOAM modulation misery

I'd like to see some details about how to determine that the ratio of 1:50 is small enough for AM:PM.

* What have people achieved in past according to the elogs©  of the measurements?

* What do we expect the effect of 1:50 to be? How much offset does this make in the MICH/PRC/SRC loops? How much offset is too much?

Recall that we are using frontal modulation with a rather small Schnupp Asymmetry...

  5502   Wed Sep 21 16:44:18 2011 KeikoUpdateIOOAM modulation mistery

AM modulation depths are found to be 50 times smaller than PM modulation depths.

m(AM,f1) ~ m(AM, f2) = 0.003 while m(PM, f1)=0.17 and m(PM, f2)=0.19.

Measured values;

* DC power = 5.2V which is assumed to be 0.74mW according to the PDA255 manual.

*AM_f1 and AM_f2 power = -55.9 dBm = 2.5 * 10^(-9) W.

P92101381.jpg

AM f2 power is assumed to be the similar value of f1. I can't measure f2 (55MHz) level properly because the PD (PDA255) is 50MHz bandwidth. From the (P_SB/P_CR) = (m/2) ^2 relation where P_SB and P_CR are the sideband and carrier power, respectively, I estimated the rough the AM modulation depths. Although DC power include the AM SB powers, I assumed that SB powers are enough small and the DC power can be considered as the carrier power, P_CR. The resulting modulation depth is about 0.003.

On the other hand, from the OSA, today's PM mod depths are 0.17 and 0.19 for f1 and f2, respectively. Please note that these numbers contains (small) AM sidebands components too. Comparing with the PM and AM sideband depths, AM sidebands seems to be enough small.

Quote:

Keiko, Suresh

AM modulations are still there ... the mechanical design for the stages, RF cables, and connections are not good and affecting the alignment. 

 

Attachment 1: P9210138.JPG
P9210138.JPG
  5501   Wed Sep 21 16:31:28 2011 PaulUpdateSUSITMY and SRM actuator response functions

 I divided the open loop transfer functions by the filter response and the sensor responses (previously measured calibration factors) to leave just the actuator responses. I've attached the actuator responses plotted in radians/count and phase over frequency.

Next step: fit the actuator response with poles and zeros.

EDIT: I divided by the wrong filter function earlier - the plots there now are divided by the correct filter function

Attachment 1: ITMY_PITCH_actuator_response.png
ITMY_PITCH_actuator_response.png
Attachment 2: ITMY_YAW_actuator_response.png
ITMY_YAW_actuator_response.png
Attachment 3: SRM_PITCH_actuator_response.png
SRM_PITCH_actuator_response.png
Attachment 4: SRM_YAW_actuator_response.png
SRM_YAW_actuator_response.png
  5500   Wed Sep 21 16:22:14 2011 ranaUpdateSUSSummary screen

The SUS SUMMARY screen is now fully activated. You should keep it open at all times as a diagnostic of the suspensions.

No matter how cool you think you are, you are probably doing something bad when trying to lock, measure any loop gains, set matrices, etc. Use the screen.

 

This is the link to the automatic snapshot of the SUS SUMMARY screen. You can use it to check the Suspensions status with your jPhone.

Auto SUS SUMMARY Snapshot

When the values go yellow its near the bad level. When its red, it means the optic is misaligned or not damped or has the wrong gain, etc.

So don't ignore it Steve! If you think the thresholds are set too low then change them to the appropriate level with the scripts is SUS/

  5499   Wed Sep 21 14:44:25 2011 PaulUpdateSUSITMY and SRM open loop transfer functions

 

 Here are the open loop transfer functions for ITMY and SRM. The various settings for the OLTFs were as follows:

Oplev filter used for all OLTFs: 300^2:0

Gains for oplev servos (for each OLTF only the 1 servo for the measured TF was on. They are all set back to 0 now):

SRM yaw gain = 1

SRM pitch gain = -1

ITMY yaw gain = -1

ITMY pitch gain = 1

measurement band = 0.2Hz to 200Hz

points = 33

swept sine magnitude envelope: amp = 2 for f > 60Hz, amp = 0.1 for f < 60Hz

Measurement points were from e.g. C1-SUS-ITMY-OLPIT-IN2 to C1-SUS-ITMY-OLPIT-IN1 to give a TF of -(loop gain).

Next step is to divide this through by the sensor reponse (i.e. the calibration factor measured earlier) and the filter response to get just the actuator response. 

 

Attachment 1: ITMY_SRM_oplev_OLTFs.png
ITMY_SRM_oplev_OLTFs.png
  5498   Wed Sep 21 14:28:25 2011 KojiSummaryLSCThe LSC code/screen modification for LSC LOCKINs

The LSC code has been modified

- The code was modified, compiled, and installed.

- The code is now running. FB was restarted to deal with the change of the channel names.

- Now we have LOCKIN1, 2, and 3. This required the change of the names from C1:LSC-LOCKIN_.... to C1:LSC-LOCKIN1_...

 

- The LSC screen has also modified. It has three lockins on the screen.

- The corresponding matrix screens have been modified/created and linked from the main screen.

- I need to make the screens more cool but the locking team can start to use those lockins.

  5497   Wed Sep 21 11:35:07 2011 steveUpdateVACRGA scan

RGA scan with maglev pumping speed at day 14 of the pump down.

The larger inserted box contains the tuning parameters of the SRS  200 amu RGA

Attachment 1: rgascanpd71m14d.jpg
rgascanpd71m14d.jpg
  5496   Wed Sep 21 09:10:15 2011 PaulUpdateSUSITMY and SRM oplev calibrations - measured and estimated

Quote:

I found that some of the Optical Lever Servos were ON today and injecting nonsense into the interferometer optics. I have set all of the gains = 0 to save us more headaches.

Please leave them OFF until we review the servo and noise characterization results in the elog.

 I had previously set the gains to zero, see the first line of my entry on Monday 5468. I should have the servo and noise characterisation done today for these oplevs today, so we can review it soon.

  5495   Wed Sep 21 02:49:39 2011 KeikoSummaryLSCLSC matrices

I created 3 kinds of LSC matrices, PRMI condition with carrier resonant in PRC, PRMI condition with SB resonant in PRC, and DRMI with SB resonant in PRC. The matrices are with AS55 and REFL11 which are used for locking right now. The signal numbers are written in log10, and the dem phases are shown in degrees.

From CR reso PRMI to SB reso PRMI, demodulation phases change  ----

 

PRMI - Carrier resonant in PRC

 

            PRCL      MICH  SRCL

REFL11 7.7079 2.9578 0
REFL33 5.2054 3.2161 0
REFL55 7.7082 2.9584 0
REFL165 3.9294 2.5317 0
AS11 1.0324 3.5589 0
AS33 1.0286 1.6028 0
AS55 1.1708 4.2588 0
AS165 1.1241 0.9352 0
POP11 2.8015 -1.3331 0
POP33 0.2989 -1.6806 0
POP55 2.8017 -0.6493 0
POP165 -0.9769 -2.3708 0
POX11 3.7954 -0.3363 0
POX33 1.293 -0.7058 0
POX55 3.796 0.355 0
POX165 0.0187 -1.3837 0
       
Dem Phase      
REFL11 3 179 0
REFL33 165 -172 0
REFL55 13 170 0
REFL165 86 177 0
AS11 -32 73 0
AS33 176 -72 0
AS55 -41 12 0
AS165 -7 146 0
POP11 -11 -116 0
POP33 124 147 0
POP55 -54 -146 0
POP165 -117 -25 0
POX11 -87 15 0
POX33 -105 -80 0
POX55 -76 16 0
POX165 180 -91 0

PRMI - SB resonant in PRC

SB reso PRMI    
  PRCL MICH SRCL
REFL11 7.6809 5.2777 0
REFL33 5.2465 3.1565 0
REFL55 7.2937 5.589 0
REFL165 4.3892 2.6857 0
AS11 1.3123 3.545 0
AS33 0.9331 1.6022 0
AS55 1.7425 4.0514 0
AS165 1.5838 1.1344 0
POP11 2.7745 0.3791 0
POP33 0.3401 -1.7392 0
POP55 2.3872 0.6904 0
POP165 -0.5171 -2.2279 0
POX11 3.7684 1.3574 0
POX33 1.3341 -0.7664 0
POX55 3.3815 1.6688 0
POX165 0.4785 -1.2163 0
       
Dem Phase
     
REFL11 155 -115 0
REFL33 -8 3 0
REFL55 91 -178 0
REFL165 -62 28 0
AS11 109 62 0
AS33 -39 99 0
AS55 13 -38 0
AS165 -155 168 0
POP11 141 -128 0
POP33 -48 -38 0
POP55 24 115 0
POP165 95 -176 0
POX11 65 155 0
POX33 83 95 0
POX55 2 92 0
POX165 32 123 0

DRMI - SB resonant in PRC

REFL11 7.6811 5.0417 4.2237 
REFL33 5.2751 4.1144 3.7766
REFL55 7.2345 7.0288 6.6801
REFL165 4.3337 4.1266 3.7775
AS11 1.1209 3.512 0.9248
AS33 0.9159 1.6323 0.7971
AS55 2.6425 5.3915 2.5519
AS165 2.6423 2.4881 2.3272
POP11 2.7747 0.1435 -0.6846
POP33 0.3687 -0.7849 -1.122
POP55 2.3244 2.1302 1.7815
POP165 -0.5833 -0.8 -1.1548
POX11 3.7676 3.261 0.8086
POX33 1.3896 0.2372 0.2333
POX55 3.4619 3.0097 3.1326
POX165 0.782 0.6668 0.4357
                        
Dem Phase
     
REFL11 154 -16 4
REFL33 -5 12 51
REFL55 129 -166 -123
REFL165 -23 40 83
AS11 132 79 69
AS33 -92 -127 -83
AS55 -33 -55 -5
AS165 154 179 -144
POP11 141 -29 -9
POP33 -46 -27 12
POP55 62 127 170
POP165 135 -161 -117
POX11 64 -102 -83
POX33 85 143 118
POX55 57 103 124
POX165 99 155 -164

 

 

  5494   Wed Sep 21 00:37:01 2011 ranaUpdateSUSITMY and SRM oplev calibrations - measured and estimated

I found that some of the Optical Lever Servos were ON today and injecting nonsense into the interferometer optics. I have set all of the gains = 0 to save us more headaches.

Please leave them OFF until we review the servo and noise characterization results in the elog.

  5493   Wed Sep 21 00:34:29 2011 ranaUpdateSUSSUS diag stuff... just so I remember what I'm doing

ETMX was ringing up when it was mis-aligned for Y arm locking. I restored the input matrix to something more diagonal and its now damping again. Needs more work before we can use the calculated matrix.

  5492   Tue Sep 20 23:59:53 2011 KojiSummaryLSCPlan to update the LSC code for multiple lock-ins

DRMI team needs to use at least three lockins on LSC

  • Increase the number of the lockin matrix  done
  • Duplicate lockin modules in the LSC code  done
  • modify the main LSC screen done
  • modify the lockin screen done
  • modify the lockin matrix screen done
  5491   Tue Sep 20 23:01:37 2011 KeikoUpdateIOOAM modulation mistery

Keiko, Suresh

AM modulations are still there ... the mechanical design for the stages, RF cables, and connections are not good and affecting the alignment.

I write the activity in the time series this time - Because we suspect the slight EOM misalignment to the beam produces the unwanted AM sidebands, we tried to align the EOM as much as possible. First I aligned the EOM tilt aligner so that the maximum power goes through. I found that about 5% power was dumped by EOM. After adjusting the alignment, the AM modulation seemed be much better and stable, however, it came up after about 20 mins. They grew up up to about -40dBm, while the noise floor is -60 dBm (when AM is minimised, with DC power of 8V by PDA225 photodetector).

We changed the EOM stage (below the tilt aligner) from a small plate to a large plate, so that the EOM base can be more stable. The EOM stands on the pile of several black plate. There was a gap below the tilt aligner because of a small plate.  So we swapped the small plate to large plate to eliminate the springly gap. However it didn't make any difference - it is the current status and there is still AM modulations right now.

During above activities, we leaned that the main cause of the EOM misalignment may be the RF cables and the resonator box connected to the EOM. They are connected to the EOM by an SMA adaptor, not any soft cables. It is very likely applying some  torc force to the EOM box. The resonator box is almost hunging from the EOM case and just your slight touch changes EOM alinment quite a bit and AM mod becomes large. 

I will replace the SMA connector between the resonator box and EOM to be a soft cable, so that the box doesn't hung from EOM tomorrow. Also, I will measure the AM mod depth so that we compare with the PM mod depth.

 

Quote:

 Keiko, Anamaria

We started to investigate the AM modulation mistery again. Checking just after the EOM, there are AM modulation about -45dBm. Even if we adjust the HWP just before the EOM, AM components grow up in 5 mins. This is the same situation as before. Only the difference from before is that we don't have PBS and HWP between the EOM and the monitor PD. So we have a simpler setup this time.

We will try to align the pockells cell alignment tomorrow daytime, as it may be a problem when the crystal and the beam are not well parallel. This adjustment has been done before and it didn't improve AM level at that time.

 

  5490   Tue Sep 20 21:13:39 2011 SureshUpdateIOOMC aligned and PSL beam into MC readjusted

This morning after Kiwamu maximised the PSL beam coupling into the MC we noticed that the MC2 face camera showed the spot position had moved away from the center by about a diameter.  So I checked the beam spot positions with MCASS and indeed found that the spot on MC2 had moved to about 6mm away from the center in yaw and about 3mm in pitch.  I adjusted the MC2 (and only MC2) to recenter the spots on all the three mirrors.  The new spot positions are given below

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    1.3337   -0.2660    0.6641   -1.0973    0.0468   -1.7130

The PSL beam into MC has been readjusted for maximal coupling into MC.

 

  5489   Tue Sep 20 20:58:35 2011 AnamariaConfigurationLSCNew AP Table Drawing

As promised, I have made a final AP table drawing, including the MC camera relocation changes by Kiwamu. I have posted it in the wiki on the tables list, and on the AP table page I've attached the inkscape .svg I used to make it, if someone needs to do small modifications.

Attached is a pdf version of it.

Big changes:

1) REFL beam has been split into 4, to go in equal powers and equal beam size to the now 4 REFL RFPDs, 11, 33, 55 and 165. A lens had to be added for REFL165 because it's a 1mm PD instead of 2mm like the other 3.

2) MC camera has moved.

3) I've cleaned up most of the random components on the table, put them away, and tidied up the cabling.

 

Attachment 1: APtableSep20th.pdf
APtableSep20th.pdf
  5488   Tue Sep 20 19:00:49 2011 PaulUpdateSUSITMY and SRM oplev calibrations - measured and estimated

 

Kiwamu noticed that the 1/L in the counts per radian should have just been L, which accounts for most of the discrepancy. We checked the input filters on the OSEMs, and they have 10dB of gain at DC. Accounting for this, estimates on the order of 20urad/count, which is much more reasonable!

  5487   Tue Sep 20 18:03:45 2011 PaulUpdateSUSITMY and SRM oplev calibrations - measured and estimated

The measured calibration factors for the oplevs are as follows:

 
SRM pitch: 666urad per count on channel C1-SUS-SRM-OLPIT-INMON
SRM yaw: 557urad per count on channel C1-SUS-SRM-OLYAW-INMON
 
ITMY pitch: 470urad per count on channel C1-SUS-ITMY-OLPIT-INMON
ITMY yaw: 491urad per count on channel C1-SUS-ITMY-OLYAW-INMON
 
Since I'm going to calibrate all the other oplevs with the rougher technique of estimating the angle from the OSEM signals directly, I thought I would check the result of such an estimation for the oplevs I have calibrated already. My method was as follows:
 
dA = change in angle
dx = change in OSEM flag position
dV = change in OSEM PD voltage
dC = change in OSEM counts
D = optic diameter
L = distance between OSEMs = D/sqrt(2)-0.002m = 0.052m
dV/dx = OSEMs volts per meter flag position change = 1700 V/m
dC/dV = OSEM counts per volt = 2^16/40 = 65536/40 counts/V
 
counts per radian = dC/dA = dV/dx  x   dC/dV   x  1/L = 1700*65536/40/0.052 = 5.3564x10^7 counts/rad
 
radians per count = dA/dC = 1.867x10^-8, or 0.019 urad/count
 
This is around a factor of 1000 smaller than what I measured earlier, reported in entry 5468. I guess this might be an issue with the whitening filter on the OSEMs, but my initial feeling was that this was only a factor of a few. If anyone can see a big obvious mistake in my above calculations please let me know!
 
 
  5486   Tue Sep 20 17:45:30 2011 kiwamuUpdateCDSdaqd is restarting by hisself ?

[Jenne / Kiwamu]

 Fb was sick. Dataviewer and Fourier Tools didn't work for a while.

After 10 minutes later they became healthy again. No idea what exactly was going on.

One thing we found was that : during the sickness of fb, it looks like daqd was restarting by hisself. Is this normal ??

Here is the bottom sentences of restart.log. Apparently daqd was rebooting although we didn't command to do so.

  daqd_start Tue Sep 20 02:41:17 PDT 2011
  daqd_start Tue Sep 20 13:18:12 PDT 2011
  daqd_start Tue Sep 20 17:33:00 PDT 2011

  5485   Tue Sep 20 16:45:09 2011 JenneUpdateSUSSUS diag stuff... just so I remember what I'm doing

Has the Q been checked?  Still in progress...

Optic POS PIT YAW SIDE
ITMX  done  done done done
ITMY  done  done  fine??  done
ETMX  done  done  done  done
ETMY  done  done  done  done
BS  done  done  done  done
PRM done done done done
SRM done done done done
MC1        
MC2        
MC3        

 So, update as of 6:17pm:  I have tuned the damping gains for all IFO optics.  Everything is good, except for ITMY Yaw.  It's probably fine, the optic damps okay, but it doesn't look like a nice clean ringdown.  I haven't taken the time to go back and look at it again.

I have to go to a dinner, but later (probably in the morning, so I don't disturb evening locking) I'll check the MC Qs.

  5484   Tue Sep 20 16:38:25 2011 KeikoUpdateIOOSmall modulation depth

Resonator box and the modulations are back now. But the modulation depth seems to be a bit smaller than yesterday, looking at the optical spectrum analyser.

 

Quote:

 Modulation resonator box is removed and the modulation depth is small right now.

I have broke the BNC connector on the modulation resonator box. The connector was attached by the screw inside very loosely and when we connect and disconnect the BNC cables from outside, extra force was applied to the cable inside and it was broke. It is being fix by Kiwamu and will be back in a bit

 

  5483   Tue Sep 20 16:31:24 2011 KeikoUpdateIOOSmall modulation depth

 Modulation resonator box is removed and the modulation depth is small right now.

I have broke the BNC connector on the modulation resonator box. The connector was attached by the screw inside very loosely and when we connect and disconnect the BNC cables from outside, extra force was applied to the cable inside and it was broke. It is being fix by Kiwamu and will be back in a bit.

 

 

 

 

  5482   Tue Sep 20 15:54:42 2011 kiwamuUpdateCamerasMC refl camera is available

[Suresh / Kiwamu]

 The MC REFL camera is now available. The camera name is "MCR" and you can call it from the videoswitch script.

 

(what we did)

 + repositioned and aligned the MCR camera.

 + checked the MCR camera.

  => found the camera view shows a negative image (i.e. the beam spot is dark and the background is bright !!)

 + replaced the camera by a spare one.

 + modified the videoswitch script because the input channel 3 was wrongly assigned to MCR.

  MCR was correctly assigned to the input channel 18.

  5481   Tue Sep 20 15:39:57 2011 KojiUpdateSUSfree swinging test in vacuum condition

Can't we use Yuta's auto-Q adjust script?

 http://nodus.ligo.caltech.edu:8080/40m/3723

Edit by KI :

Of course we can use it but first we have to fix some pynds sentences since his script was written for the OLD pynds.

  5480   Tue Sep 20 15:23:16 2011 JenneUpdateSUSfree swinging test in vacuum condition

This is using data for the SRM from: 20 Sept 2011 03:20:00 PDT = 1000549215

You can see that there are still some funny peaks between Pit and Yaw, but I finnessed the peak-finding, and I was able to fit all of the correct peaks, and invert the matrix:

 SRM now has its new matrix, and is damping happily.

Optic The Plot Matrix Badness
SRM SRM.png                pit     yaw     pos     side    butt
UL    0.877   0.983   1.105  -0.288   1.092 
UR    1.010  -1.017   1.123  -0.145  -1.055 
LR   -0.990  -1.002   0.895  -0.091   0.848 
LL   -1.123   0.998   0.877  -0.234  -1.006 
SD    0.089   0.064   3.752   1.000  -0.009
 4.4076

 

 

  5479   Tue Sep 20 14:53:13 2011 JenneUpdateSUSJenne's Scripts started

Quote:

Quote:

Quote:

I followed Jenne's instructions, ran the matrix filler script and then set the optics to freeswing. Someone has to burt resture and damp them in the morning.

 Thanks!  I'll give them a little more time, then restore things.

 I began restoring the optics at ~9:30am, so I have a full 6 hours of data, in case I need that much to separate the Pos/Side modes on some of the optics.  They are all damping again with their original matricies.

 So, clearly this was a kind of dumb idea.  There is nothing mechanical going on between our sensor inputs and our Pit/Pos/Yaw/Side DoF filter banks.  It's just math.  On the other hand, we now have a 3rd set of in-vac free swinging data, so I can (after all the suspensions are working) have a look at the drift in matrix elements over time.

In other news, after some meditation, and fitzing with DoF gain values, all of the IFO optics except for SRM now have their new input matricies, and are damping pretty nicely.  I need to go through and do an "eyeball" check to make sure that everything has a Q of ~5ish.  So far, I've kicked the optics, and watched that they damped fairly quickly, but I don't have a guesstimate of the Q's for each optic, for each DoF.

So, still to do:

Use another set of data and invert the SRM matrix DONE

Plug in the MC matricies, make sure they're okay. DONE

Check the Q's for all optics, all DoFs. 

  5478   Tue Sep 20 13:57:44 2011 kiwamuUpdateIOOincident beam to MC aligned

Since the MC wasn't able to capture the 00 mode in this morning I aligned the incident beam going to MC.

As a result C1:IOO-RFPD_DCMON went down to 0.6. However the beam on IPPOS is almost falling off from the QPD.

  5477   Tue Sep 20 09:44:44 2011 JenneUpdateSUSJenne's Scripts started

Quote:

Quote:

I followed Jenne's instructions, ran the matrix filler script and then set the optics to freeswing. Someone has to burt resture and damp them in the morning.

 Thanks!  I'll give them a little more time, then restore things.

 I began restoring the optics at ~9:30am, so I have a full 6 hours of data, in case I need that much to separate the Pos/Side modes on some of the optics.  They are all damping again with their original matricies.

  5476   Tue Sep 20 04:12:26 2011 JenneUpdateSUSJenne's Scripts started

Quote:

I followed Jenne's instructions, ran the matrix filler script and then set the optics to freeswing. Someone has to burt resture and damp them in the morning.

 Thanks!  I'll give them a little more time, then restore things.

  5475   Tue Sep 20 03:12:14 2011 AnamariaUpdateSUSJenne's Scripts started

I followed Jenne's instructions, ran the matrix filler script and then set the optics to freeswing. Someone has to burt resture and damp them in the morning.

  5474   Tue Sep 20 03:02:23 2011 KeikoUpdateLSClocking activity tonight

 Keiko, Anamaria, Koji

We were not able to establish the stable DRMI tonight. We could lock MICH and PRCL quite OK, and lock the three degrees of freedom at somewhere strange for several seconds quite easily, but the proper DRMI lock was not obtained.

When MICH and PRC are locked to the carrier, REFL DC PD reading dropps from ~3000 counts to 2600~2700 counts as REFL beam is absorbed to PRC. We'll try to lock PRC to sidebands - but flipping gain sign didn't work today, although it worked a few days ago. 

POP beam (monitor) is useful to align PRM.

  5473   Tue Sep 20 02:21:10 2011 KojiUpdateLSCLSC MEDM screen cleaning up

I have made some cleaning up of the LSC-related MEDM screens.

- LSC overview screen: ADC OVFL and WFAA indicators are now correctly matched to it associated PD signals.

- Whitening screens now have the correct indication of the associated PD signals.

- LSC Ctrl screen, which is invoked from the overview screen by clicking the servo filters, now has the switches of the servo filters.

- LSC tab of the sitemap was cleaned up by removing the broken links.

  5472   Mon Sep 19 23:19:40 2011 KeikoUpdateIOOAM modulation mistery

 Keiko, Anamaria

We started to investigate the AM modulation mistery again. Checking just after the EOM, there are AM modulation about -45dBm. Even if we adjust the HWP just before the EOM, AM components grow up in 5 mins. This is the same situation as before. Only the difference from before is that we don't have PBS and HWP between the EOM and the monitor PD. So we have a simpler setup this time.

We will try to align the pockells cell alignment tomorrow daytime, as it may be a problem when the crystal and the beam are not well parallel. This adjustment has been done before and it didn't improve AM level at that time.

  5471   Mon Sep 19 22:47:44 2011 JenneUpdateSUSSUS diag stuff... just so I remember what I'm doing

 The last person out tonight should run the following scripts:

In Matlab: 

/opt/rtcds/caltech/c1/scripts/SUS/peakFit/writeMultiSUSinmat.m

In command line:

/opt/rtcds/caltech/c1/scripts/SUS/freeswing all

 

Then in the morning, someone should do a BURT restore to early today (to get the default matricies back), and also restore the watchdogs.

Thanks!
 

  5470   Mon Sep 19 21:19:25 2011 KatrinUpdateGreen LockingBroadband photodiode characterization

Another Hamasutu S3399 photodiode was tested with the electronic circuit as described in LIGO-D-1002969-v.

RF transimpedance is 1k although the DC transimpedance is 2k.

The noise level is 25pA/sqrt(Hz) which corresponds to a dark current of 1.9mA or 1.7mA in the independent measurement.

At all frequencies the noise is larger compared to Koji's measurement (see labbook page 4778).

 


In file idet_S3399.pdf the first point is not within its error bars on the fitted curved. This point corresponds to the dark noise measurement

I made this measurement again. Now it is on the fitted curve. In the previous measurement I pushed the save button a bit too early. The

averaging process has not been ready while I pushed the 'save'  button.

Dark current is 1.05mA and noise is lower than in the previous measurement.

New file are the XXX_v2.pdf files

current_noise_S3399_v2.pdf

 idet_S3399_v2.pdf

 


 idet_S3399.pdf

 current_noise_S3399.pdf

S3399_response.pdf

  5468   Mon Sep 19 20:56:36 2011 PaulSummarySUSRemaining SRM and ITMY OSEMs calibrations

 

I've now taken data for the pitch and yaw calibrations for the OSEMs of SRM and ITMY. Until such time as I know what the calibrated oplev noise spectra are like, I'm leaving the servo gains at zero.

I estimate the length of the lever arm from SRM to measurement position to be 3.06m, and the length of the lever arm from the ITMY to the measurement position to be 3.13m.

From the fits shown on the attached plots, this gives the following calibration factors for the SRM and ITMY OSEMs pitch and yaw counts (i.e. counts from channels such as SUS-ITMY_ULSEN_SW2 multiplied by a matrix of 1s and -1s) to pitch and yaw angle:

 

SRM PITCH: 1 OSEMs pitch count = 11.74 microradians

SRM YAW: 1 OSEMs yaw count = 12.73 microradians

 

ITMY PITCH: 1 OSEMs pitch count = 13.18 microradians

ITMY YAW: 1 OSEMs yaw count = 13.52 microradians

 

Next step is to do some DC offsets with the oplev paths back in place to get the final calibration between OSEMs counts and oplev counts, thus finally getting a conversion factor from oplev counts to radians.

I noticed while taking these measurements that the DC offsets I put on ITMY caused around 5 times larger change in angle than those on the SRM. The different path length is not enough to account for this, so I propose that the actuation is working differently for the two. I guess this should be taken into account when designing the output matrices (unless the control is passed through a different output matrix than the DC offsets?). I'll quantify the difference shortly, and write a conversion factor between output alignment count (e.g. SUS-ITMY_PIT_COMM) and angle.

 

 

Attachment 1: SRM_PITCH_calib_curve.png
SRM_PITCH_calib_curve.png
Attachment 2: SRM_YAW_calib_curve.png
SRM_YAW_calib_curve.png
Attachment 3: ITMY_PITCH_calib_curve.png
ITMY_PITCH_calib_curve.png
Attachment 4: ITMY_YAW_calib_curve.png
ITMY_YAW_calib_curve.png
ELOG V3.1.3-